Circular No. 5630

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1992bc IN ESO 300-G9 M. Hamuy, Cerro Tololo Interamerican Observatory; and J. Maza,

University of Chile, report the discovery by R. Antezana (University of Chile) of a supernova located ∼ 15′′ east and 5′′ north of the nucleus of the spiral galaxy ESO 300-G9 (α = 3^{h}03^{m}22^{s}.2, δ = −39^{o}45′15′′, equinox 1950.0). The supernova, of m_{pg} ∼ 15, was found on a 20-min unfiltered IIa-O plate taken by C. Muena and G. Valladares with the CTIO Curtis Schmidt telescope on Oct. 4.364 UT. A spectrogram (range 320.0-750.0 nm), obtained by Maza on Oct. 6.1 with the CTIO 4-m telescope, shows the characteristic Si II 635.5-nm feature of type-Ia supernovae. The minimum of the Si II absorption yields an expansion velocity of ∼ 13 000 km/s with respect to the galaxy’s rest frame. This spectrum shows resemblance to that of the type-Ia SN 1989B, 7 days before maximum light. The spectrum of the galaxy’s nucleus yields a redshift of z = 0.019.

SUPERNOVA 1978K IN NGC 1313 A. V. Filippenko, University of California at Berkeley, writes: “Careful

consideration of the available data strongly supports the suggestion of Ryder et al. (IAUC 5615) that the suspected nova in NGC 1313 (IAUC 4950) is actually a supernova. The x-ray luminosity exceeds 10^{40} erg/s, and the nonthermal radio luminosity is over 200 times that of Cas A, the brightest Galactic supernova remnant. It is probable that the progenitor was a very massive star having a dense, slow stellar wind with a mass-loss rate exceeding 10^{−4} solar masses per year. The observed properties resemble those of SN 1986J in NGC 891, and perhaps of SN 1961V in NGC 1058. Based on its measured brightness, SN 1978K is extinguished by roughly 4-6 mag at optical wavelengths, possibly due to the dense circumstellar gas. Further study of this object, which will help characterize the growing subclass of radio- and x-ray-luminous type-II supernovae, is encouraged.”

SUPERNOVA 1992ba IN NGC 2082 A. C. Gilmore reports the following position of SN 1992ba, measured by

P. M. Kilmartin from films taken by Gilmore with the 0.6-m f/14 Cassegrain telescope and 0.15-m f/15 astrographs at Mt. John University Observatory on Oct. 4.62 UT: α = 5^{h}41^{m}31^{s}.88, δ = −64^{o}19′21′′.0 (equinox 1950.0, uncertainty ∼ 1′′ in each coordinate).

1992 October 6 Daniel W. E. Green